OpenStudy (anonymous):

Please Help o 1. If x is a positive integer and the units digit of x2 is 9 and the units digit of (x+1)2 is 4, what is the units digit of (x+2)2?

hartnn (hartnn):

@LuizaMiller last digit if x can be either 3 or 7 because 3^2=9, 7^2=49 if its 3, then last digit of x+1 = 4, and its square = 16, so last digit of (x+1)^2 will be 6, but it is not! so, take last digit as 7 (x+1) = 8, last digit of (x+1)^2 = 4 and (x+2) = 9 , and last (unit) digit will be 1 (as 9^2=81)